The balance sheet provides a snapshot of the financial condition of a company. Investors and analysts use the information given on the balance sheet and other financial statements to make several interpretations regarding the companys financial condition and performance Fuzzy Button Clothing Company is a hypothetical company. Suppose it has the following balance sheet items reported at the end of its first year of operation. For the second year, some parts are still incomplete. Use the information given to complete the balance sheet. Fuzzy Button Clothing Company Balance Sheet for Year Ending December 31 (Millions of Dollars) Year 2 Year 1 Year 2 Year 1 Assets Liabilities and equity Current liabilities: Current assets: Cash and equivalents Accounts receivable Inventories $1,845Accounts payable $0 $0 844 2,475 $5,625 675 1,980 $4,500 Accruals 664 625 $625 1,875 $2,500 Notes payable Total current assets Total current liabilities Net fixed assets: Long-term debt 2,344 $3,125 Net plant and equipment $5,500 Total debt Common equity: 4,875 2,625 $7,500 $10,000 Common stock 6,094 Retained earnings Total common equity Total liabilities and equity $9,375 $12,500 Total assets $12,500 $10,000 Given the information in the preceding balance sheet-and assuming that Fuzzy Button Clothing Company has 50 million shares of common stock outstanding-read each of the following statements, then identify the selection that best interprets the information conveyed by the balance sheet.
Answer #1

Total current liabilities in year 2 = 117 + 664 = 781
Retained earnings in year 2 = Total equity - Common stock = 9375 - 6094 = 3281

Net plant and equipments = Total assets - Total current assets = 12500 - 5625 = 6875
Cash and Cash equivalent = 5635 - 844 -2475 = 2306

Statement 1 is false because option c is correct.
The accrual has increase drom 0 to 117

Statement 2 is true beacuse option c is correct.
Retained earnings has increased from year 1 to year 2

Statement 3 is true beacuse option b is correct
Book value per share = Equity value/No of shares outstanding.9375/50 = 187.50

Last question Option c is correct.
As per GAAP the debt should be arranged in increasing order of liquidity.

Please Discuss in case of doubts
Best of Luck. God Bless
Please rate Well